# Difference-in-Discontinuities: Estimation, Inference and Validity Tests

### v1

@@ 5.2 Time-varying DGPs
Yes confounder
RDD (Robust) 1.040 1.040 0 0.289 0.074 0.215 0.348
DiDC (RDD of $\Delta$s) 0.016 0.067 0.906 0.290 0.074 0.195 0.337
DiDC ($\Delta$ of RDDs) 0.014 0.066 0.914 0.290 0.074 0.214 0.348
DiD -1.568 1.568 0 0.384 0.098 - -
OLS of $\Delta$s 0.684 0.684 0 0.244 0.363 - -
No confounder
RDD (Robust) 0.040 0.069 0.899 0.291 0.074 0.215 0.348
DiDC (RDD of $\Delta$s) 0.016 0.067 0.905 0.290 0.074 0.195 0.336
DiDC ($\Delta$ of RDDs) 0.014 0.066 0.914 0.290 0.074 0.214 0.348
DiD -1.568 1.568 0 0.384 0.098 - -
OLS of $\Delta$s 0.684 0.684 0 0.244 0.363 - -
Note: Simulations based on 10,000 Monte Carlo experiments with a sample size $n=500$. RDD is the non-bias corrected RD estimator from CCT (2014), DiD is the outcome regression DiD estimator from Sant’Anna and Zhao (2020), DiDC is the estimator proposed in this paper. “Av. Bias”, “RMSE”, “Cover”, “CIL’ and "Av. SE, stand for the average simulated bias, simulated root mean-squared errors, 95% coverage probability, 95% confidence interval length and average standard error, respectively. "Bandwidths" $h_{n}$ and $b_{n}$ report the plug-in bandwidths for point and bias estimation, respectively.
This table shows that in the scenario without confounding factors and time-varying DGPs, the DiD cannot be used due to the lack of parallel trends. However, the RDD can be employed, but it shows a higher bias compared to the differences-in-discontinuity approach because the GDPs, although different, are similar enough.
With confounding factors, both the RDD and DiD estimators are unsuitable. RDD remains problematic due to the confounding effects as it does not account for pre-existing discontinuities. DiD is also unsuitable because no parallel trends exist in this context. The differences-in-discontinuities method is the only approach that correctly estimates the treatment effect.
@@ 6 Empirical Illustration
We demonstrate the application of our estimator by revisiting Grembi et al. (2016), which uses data from Italian municipalities to analyze the impact of fiscal rules on municipal deficit by exploiting a government fiscal rule relaxation as a natural experiment. Our objective is to implement the DiDC approach and compare our results with the original findings. Additionally, we conduct the validity test from Section 4.1 to assess the suitability of using DiDC in this context.
The dataset comprises data from Italian municipalities, focusing on the period surrounding the government’s relaxation of fiscal rules in 2001. Municipalities with fewer than 5,000 inhabitants experienced a relaxation of fiscal rules and are defined as treatment, while those with more than 5,000 inhabitants served as controls.
We reframe the data into a more manageable 2x2 design, simplifying the estimation. This transformation involves computing the differences in outcomes before and after the policy change. Grembi et al. (2016)’s original study utilized a large panel dataset, a rectangular kernel and a polynomial of degree one as in the model below:
$Y_{it}=$ $\delta_{0}+\delta_{1}(X_{i}-c)+S_{i}(\gamma_{0}+\gamma_{1}(X_{i}-c))$ $T_{i}[\alpha_{0}+\alpha_{1}(X_{i}-c)+S_{i}(\beta_{0}+\beta_{1}(X_{i}-c))]+\varepsilon_{it}.$
@@
For our implementation, we employ the methodology described in Section 3. Table 4 compares the results from both our DiDC methods (RDD of the differences and the differences of RDDs) to Grembi et al. (2016)’s estimations. It also reports bandwidths for point estimation and number of observations.
@@
Diff-in-Disc (RDD of $\Delta$)
23.100
44.508
-21.645
Bandwidth 619 600 420
Observations 1245 1028 1030
Diff-in-Disc ($\Delta$ of RDDs)
26.864
56.043
-18.876
Bandwidth 995 809 819
Observations 1456 1161 1164
We conduct the test from Section 4.1 to evaluate the assumption of time-invariant confounding effects. We use data from 1997 to 2000 to perform "stacked" RDD regressions. The results are reported in Table 5.
Table 5:
Dependent variable:
imposte
Z:$T_{1997}$-0.040
Z:$T_{1998}$ 0.265∗∗∗
Z:$T_{1999}$ 0.176∗∗
Z:$T_{2000}$ 0.054
Z:$T_{1997}$:$D_{1997}$ 0.280∗∗∗
Z:$T_{1998}$:$D_{1998}$-0.369∗∗∗
Z:$T_{1999}$:$D_{1999}$-0.213∗∗
Z:$T_{2000}$:$D_{2000}$ 0.038
Constant 121.487∗∗∗
Observations 4,810
R2 0.059
Adjusted R2 0.048
Residual Std. Error 2.137 (df = 707)
F Statistic 5.518∗∗∗ (df = 8; 707)
Note: p$<$0.1; ∗∗p$<$0.05; ∗∗∗p$<$0.01
The hypothesis test for $Z:T_{1997}:D_{1997}=Z:T_{1998}:D_{1998}=Z:T_{1999}:D_{1999}=Z:T_{2000}:D_{2000}$ indicates a rejection of the null hypothesis, suggesting that the coefficients are not constant over time, as reported in Table 6:
Table 6:
Statistic N Mean St. Dev. Min Max
Res.Df 2 708.500 2.121 707 710
RSS 2 3,298.776 98.196 3,229.341 3,368.211
Df 1 3.000 3 3
Sum of Sq 1 138.871 138.871 138.871
F 1 10.134 10.134 10.134
Pr(>F) 1 0.00000 0.00000 0.00000
@@
In summary, our estimated treatment effects are similar in magnitude to the difference-in-discontinuities approach utilized in Grembi et al. (2016) study. Notably, our difference of RDDs approach yields smaller confidence intervals, suggesting it to be the most powerful approach. However, our validity test reveals that confounding effects vary over time, so the use of DiDC in this setting appears to be leading to biased estimates. We also replicate our estimates using Ludwig and Miller (2007)’s bandwidths with similar results. These estimates are shown in Table E.1 in Appendix E.

### v2

@@ 6.2 Time-varying Functional Forms
RDD (Robust) 1.039 1.039 0 0.161 0.221 0.360
Diff-in-Disc -0.001 0.050 0.937 0.241 0.184 0.331
NP-DiD 1.766 1.766 0 0.297 - -
TWFE -1.568 1.568 0 0.272 - -
Note: Simulations based on 10,000 Monte Carlo experiments with a sample size $n=1000$. RDD is the non-bias corrected RD estimator from CCT (2014), NP-DiD is the outcome regression DiD estimator from Sant’Anna and Zhao (2020), DiDC is the estimator proposed in this paper. “Av. Bias”, “RMSE”, “Cover”, “CIL’ stand for the average simulated bias, simulated root mean-squared errors, 95% coverage probability and 95% confidence interval length, respectively. "Bandwidths" $h_{n}$ and $b_{n}$ report the plug-in bandwidths for point and bias estimation, respectively.
Table 3 shows that in the presence of changes in functional forms over time and confounding policies at the threshold, only the Diff-in-Disc approach is valid. The finite-sample bias of the standard RD estimator is close to the one presented in Table 1. The main difference in Table 3 in comparison to Table 1 is the poor finite-sample properties of DiD methods, as both the nonparametric DiD and the TWFE estimator exhibit larger finite-sample bias than the standard RD.
Table 4: Time-Varying Functional Forms without Confounding
AV.Bias RMSE Coverage CIL Bandwidths
$h_{n}$ $b_{n}$
RDD (Robust) 0.039 0.049 0.922 0.161 0.221 0.360
Diff-in-Disc -0.008 0.050 0.937 0.241 0.184 0.331
NP-DiD 1.766 1.766 0 0.297 - -
TWFE -1.568 1.568 0 0.272 - -
Note: Simulations based on 10,000 Monte Carlo experiments with a sample size $n=1000$. RDD is the non-bias corrected RD estimator from CCT (2014), NP-DiD is the outcome regression DiD estimator from Sant’Anna and Zhao (2020), DiDC is the estimator proposed in this paper. “Av. Bias”, “RMSE”, “Cover”, “CIL’ stand for the average simulated bias, simulated root mean-squared errors, 95% coverage probability and 95% confidence interval length, respectively. "Bandwidths" $h_{n}$ and $b_{n}$ report the plug-in bandwidths for point and bias estimation, respectively.
Table 4 displays the results for the case where functional forms change over time, but there is no confounding policy at the threshold. Once again, the standard RD and the Diff-in-Disc exhibit desirable finite-sample properties, with the Diff-in-Disc showing better coverage and smaller bias. However, unlike when functional forms are constant over time, when functional forms change, the standard RD method yields larger optimal bandwidths. For the DiD estimators, the results show severe bias, and the performance of the nonparametric DiD and the TWFE are similar to what we observed in Table 3. In Appendix D.2, we show that the simulation results are robust to changes in the time-varying variance.
Overall, the Monte Carlo results show several desirable properties of the Diff-in-Disc estimator. Not only it remains unbiased when the alternative approaches are not valid, but it also exhibits smaller finite-sample bias than the standard RD even in the absence of confounders. In the next section, we revisit a well-known political economy setting to analyze the estimator’s performance on a real dataset.
@@ 7 Empirical Illustration
We illustrate the use of our estimator by revisiting the empirical application in Grembi et al. (2016), which analyzes the impact of fiscal rules on Italian municipal finances by exploiting a 2001 fiscal rule relaxation as a natural experiment. The relaxation applied to municipalities with fewer than 5,000 inhabitants, which therefore form the treatment group, while municipalities above this threshold serve as controls.
Our objective is to replicate their setting using the Difference-in-Discontinuities (DiDC) estimator, compare the results to the original findings, and assess the validity of the identifying assumption using the constant-confounder test from Section 4.1.
The dataset comprises data from Italian municipalities, focusing on the period surrounding the government’s relaxation of fiscal rules in 2001. Municipalities with fewer than 5,000 inhabitants experienced a relaxation of fiscal rules and were treated, while those with more than 5,000 inhabitants served as controls.
We implement a 2×2 DiDC design, computing before–and–after differences and estimating local linear RDDs as detailed in Section 3. Grembi et al. (2016)’s original study utilized a large panel dataset, a rectangular kernel and a polynomial of degree one as in the model below:
$Y_{it}=$ $\delta_{0}+\delta_{1}(X_{it}-c)+S_{i}(\gamma_{0}+\gamma_{1}(X_{it}-c))$ $T_{i}[\alpha_{0}+\alpha_{1}(X_{it}-c)+S_{i}(\beta_{0}+\beta_{1}(X_{it}-c))]+\varepsilon_{it}.$
@@
Table 5 compares our DiDC estimates (difference of RDDs), along with the effective bandwidths and sample sizes to Grembi et al. (2016)’s estimations. Due to differences in specifications, the results must be interpreted with caution.
@@
Diff-in-Disc (RD of $\Delta$)
18.501
10.247
-5.288
Bandwidth 520 591 592
Observations 392 433 434
Diff-in-Disc ($\Delta$ of RDs)
22.142
9.210
-9.009
Bandwidth 560 475 334
Observations 415 362 264
Note: The first panel presents the estimates from Grembi et al. (2016) using the whole panel from 1999 to 2004. The second panel presents the estimates obtained from the RD of first-differences using data from 2000 and 2001. The third panel presents the estimates obtained from the difference of RDs using data from 2000 and 2001.
The results are qualitatively similar, yet, our specification is underpowered due to the smaller sample size, and thus estimates are not statistically significant.
7.1 Testing the Assumptions
We conduct the test from Section 4.1 to evaluate the assumption of time-invariant confounding effects. To test this, we estimate stacked RDs for the years 1998–2000, interacting the running variable with treatment and period dummies. Table 6 reports the interaction coefficients ($z\times D$) for the smallest bandwidth, and Table 7 summarizes the corresponding joint Wald tests for equality of these coefficients across years.
Table 6: Stacked-RD interaction coefficients by pre-period (smallest bandwidth)
Outcome $\theta_{2000}$ $\theta_{1999}$ $\theta_{1998}$
Taxes 1.582 1.373 1.248
(0.083) (0.084) (0.084)
Deficit 0.095 0.114 0.112
(0.020) (0.020) (0.021)
Fiscal Gap 1.273 1.338 1.290
(0.073) (0.074) (0.075)
Note: Estimates from weighted stacked OLS (triangular kernel). Standard errors in parentheses. “Smallest bandwidth” = minimum period-specific CCT bandwidth.
The joint test of $H_{0}:\theta_{2000}=\theta_{1999}=\theta_{1998}$ is implemented using a Wald $F$-test based on the pooled regressions. The results are shown in Table 7.
Table 7: Joint Wald test: $H_{0}:\theta_{2000}=\theta_{1999}=\theta_{1998}$ (F-stat, df, p-value)
Outcome Bandwidth F-statistic p-value
Taxes Smallest 4.108 0.017
Biggest 4.545 0.011
CCT 3.010 0.046
Deficit Smallest 0.279 0.757
Biggest 0.181 0.835
CCT 0.294 0.745
Fiscal balance Smallest 0.213 0.809
Biggest 0.141 0.869
CCT 0.182 0.833
Note:F-statistics and p-values are from the joint Wald test implemented with car::linearHypothesis (Fox et al., 2001). Bold p-values indicate rejection at 5%.
For taxes, the null hypothesis of time-invariant confounding is rejected at conventional significance levels (p $\in$ [0.01, 0.05]), indicating that the RD discontinuity varied across pre-treatment years. In contrast, the test fails to reject $H_{0}$ for both deficit and fiscal balance (p-values between 0.75 and 0.87).
@@ 7.1 Testing the Assumptions
In summary, our estimated treatment effects are similar in magnitude to the difference-in-discontinuities approach utilized in Grembi et al. (2016) study. Notably, our difference-in-RDDs approach yields smaller confidence intervals, suggesting it is the most powerful. However, our validity test indicates that confounding effects vary over time, suggesting that the use of DiDC in this setting leads to biased estimates. We also replicate our estimates using the bandwidths in Ludwig and Miller (2007), with similar results. These estimates are shown in Table E.9 in Appendix E.
We also implement the KS-type test from Section 4.2 to assess whether the shapes of the conditional mean functions remain stable across pre-treatment years. Applying this test to each outcome, we find strong evidence of time-varying functional forms for taxes and fiscal balance: on both sides of the cutoff, the bootstrap p-values for imposte and saldo are essentially zero (0.001 on both sides), indicating clear violations of the time-invariance condition. For the deficit, however, the evidence is mixed. The left-hand side yields a p-value of 0.266, consistent with time-invariant functional forms, whereas the right-hand side again produces a very small p-value (0.001), suggesting that the relationship between the running variable and the deficit outcome changed over time for municipalities above the 5,000-inhabitant threshold. Overall, these results show that the shape of the outcome–running-variable relationship is not stable across pre-treatment years for most outcomes, reinforcing the need for careful consideration when applying the method.
7.2 Partial Identification
7.2.1 Bounded Variation Assumptions
Table 8 reports the identified sets for the treatment effect on Taxes under the bounded-variation sensitivity framework. Each cell shows the lower and upper bound for the identified set $\left[\tau_{c}^{LB},\,\tau_{c}^{UB}\right]$ as a function of the sensitivity parameters $c_{1}$ and $c_{2}$. The bounds are constructed by allowing the conditional potential-outcome functions and the confounding effect to deviate from the time-invariance baseline up to $c_{1}$ and $c_{2}$, respectively (Section 5). The same grid procedure was used to produce the analogous tables for Fiscal balance (Tables E.10E.11 in Appendix E) and Deficit (Tables 12 and 13 from Section 7.2.3).
Table 8: Bounds on the Treatment Effect under Bounded Variation (Outcome: Taxes)
$c_{2}\backslash c_{1}$ 0.0 0.5 1.0 1.5 2.0 2.5 3.0 3.5 4.0 4.5 5.0
0.0 6.33 , -6.57 5.83 , -6.57 5.33 , -6.57 4.83 , -6.57 4.33 , -6.57 3.83 , -6.57 3.33 , -6.57 2.83 , -6.57 2.33 , -6.57 1.83 , -6.57 1.33 , -6.57
0.5 6.33 , -6.07 5.83 , -6.07 5.33 , -6.07 4.83 , -6.07 4.33 , -6.07 3.83 , -6.07 3.33 , -6.07 2.83 , -6.07 2.33 , -6.07 1.83 , -6.07 1.33 , -6.07
1.0 6.33 , -5.57 5.83 , -5.57 5.33 , -5.57 4.83 , -5.57 4.33 , -5.57 3.83 , -5.57 3.33 , -5.57 2.83 , -5.57 2.33 , -5.57 1.83 , -5.57 1.33 , -5.57
1.5 6.33 , -5.07 5.83 , -5.07 5.33 , -5.07 4.83 , -5.07 4.33 , -5.07 3.83 , -5.07 3.33 , -5.07 2.83 , -5.07 2.33 , -5.07 1.83 , -5.07 1.33 , -5.07
2.0 6.33 , -4.57 5.83 , -4.57 5.33 , -4.57 4.83 , -4.57 4.33 , -4.57 3.83 , -4.57 3.33 , -4.57 2.83 , -4.57 2.33 , -4.57 1.83 , -4.57 1.33 , -4.57
2.5 6.33 , -4.07 5.83 , -4.07 5.33 , -4.07 4.83 , -4.07 4.33 , -4.07 3.83 , -4.07 3.33 , -4.07 2.83 , -4.07 2.33 , -4.07 1.83 , -4.07 1.33 , -4.07
3.0 6.33 , -3.57 5.83 , -3.57 5.33 , -3.57 4.83 , -3.57 4.33 , -3.57 3.83 , -3.57 3.33 , -3.57 2.83 , -3.57 2.33 , -3.57 1.83 , -3.57 1.33 , -3.57
3.5 6.33 , -3.07 5.83 , -3.07 5.33 , -3.07 4.83 , -3.07 4.33 , -3.07 3.83 , -3.07 3.33 , -3.07 2.83 , -3.07 2.33 , -3.07 1.83 , -3.07 1.33 , -3.07
4.0 6.33 , -2.57 5.83 , -2.57 5.33 , -2.57 4.83 , -2.57 4.33 , -2.57 3.83 , -2.57 3.33 , -2.57 2.83 , -2.57 2.33 , -2.57 1.83 , -2.57 1.33 , -2.57
4.5 6.33 , -2.07 5.83 , -2.07 5.33 , -2.07 4.83 , -2.07 4.33 , -2.07 3.83 , -2.07 3.33 , -2.07 2.83 , -2.07 2.33 , -2.07 1.83 , -2.07 1.33 , -2.07
5.0 6.33 , -1.57 5.83 , -1.57 5.33 , -1.57 4.83 , -1.57 4.33 , -1.57 3.83 , -1.57 3.33 , -1.57 2.83 , -1.57 2.33 , -1.57 1.83 , -1.57 1.33 , -1.57
Note: Estimates for the identified sets for the effects on fiscal rules on taxes under Lemma 6.
7.2.2 Modularity Assumptions
Under Assumptions 7 and 8, we construct worst-case modularity bounds using $y_{min}$ and $y_{max}$ equal to the minimum and maximum observed values of each outcome in the data (Table 9). As expected in this worst-case setting, the resulting identified sets are very wide: for example, the bounds for taxes span roughly [157.59, 1203.19][-157.59,\;1203.19] for $\tau_{c}$ and [1203.19, 108.08][-1203.19,\;108.08] for $\tau_{uc}$, with similarly large intervals for the fiscal gap and deficit. Consequently, these bounds don’t provide much information. To obtain informative conclusions about the direction or magnitude of the treatment effect, researchers must impose stronger and more economically grounded restrictions.
Table 9: Identified Sets Under Modularity Assumptions
Outcome $\tau_{c}\in[\tau_{c}^{LB},\tau_{c}^{UB}]$ $\tau_{uc}\in[\tau_{uc}^{LB},\tau_{uc}^{UB}]$
Taxes [157.59, 1203.19][-157.59,\;1203.19] [1203.19, 108.08][-1203.19,\;108.08]
Fiscal Gap [738.37, 2358.18][-738.37,\;2358.18] [2358.18, 862.38][-2358.18,\;862.38]
Deficit [757.29, 1687.17][-757.29,\;1687.17] [1687.17, 758.35][-1687.17,\;758.35]
Note: Estimates for the identified sets for the effects on fiscal rules on public finance outcomes under Lemma 7.
7.2.3 Combining Bounded Variation and Modularity Assumptions
When the two approaches are combined, identification becomes substantially sharper. Together, these restrictions eliminate many of the extreme scenarios allowed by modularity alone and narrow the identified sets relative to either assumption in isolation.
The Deficit outcome clearly illustrates the value of this combination. Under worst-case modularity, the identified sets for Deficit include both large negative and positive effects, but once bounded-variation constraints are imposed, the intervals become notably narrower and remain strictly negative across all $c_{1}$ and $c_{2}$ values considered. Table 10 shows that the bounds for $\tau_{c}$ fall roughly between 27-27 and 19-19, and Table 11 similarly reports negative upper endpoints for $\tau_{uc}$. The combined assumptions, therefore, identify not only the sign of the effect but also restrict its magnitude for $\tau_{c}$ to a reasonably small range.
Table 10: Bounds for $\tau_{c}$ — Outcome: Deficit. $y_{min}=-762.5508$, $y_{max}=924.6157$.
$c_{2}\backslash c_{1}$ 0.0 0.5 1.0 1.5 2.0 2.5 3.0 3.5 4.0 4.5 5.0
0.0 19.60 , 27.86 20.10 , 27.86 20.60 , 27.86 21.10 , 27.86 21.60 , 27.86 22.10 , 27.86 22.60 , 27.86 23.10 , 27.86 23.60 , 27.86 24.10 , 27.86 24.60 , 27.86
0.5 19.60 , 27.36 20.10 , 27.36 20.60 , 27.36 21.10 , 27.36 21.60 , 27.36 22.10 , 27.36 22.60 , 27.36 23.10 , 27.36 23.60 , 27.36 24.10 , 27.36 24.60 , 27.36
1.0 19.60 , 26.86 20.10 , 26.86 20.60 , 26.86 21.10 , 26.86 21.60 , 26.86 22.10 , 26.86 22.60 , 26.86 23.10 , 26.86 23.60 , 26.86 24.10 , 26.86 24.60 , 26.86
1.5 19.60 , 26.36 20.10 , 26.36 20.60 , 26.36 21.10 , 26.36 21.60 , 26.36 22.10 , 26.36 22.60 , 26.36 23.10 , 26.36 23.60 , 26.36 24.10 , 26.36 24.60 , 26.36
2.0 19.60 , 25.86 20.10 , 25.86 20.60 , 25.86 21.10 , 25.86 21.60 , 25.86 22.10 , 25.86 22.60 , 25.86 23.10 , 25.86 23.60 , 25.86 24.10 , 25.86 24.60 , 25.86
2.5 19.60 , 25.36 20.10 , 25.36 20.60 , 25.36 21.10 , 25.36 21.60 , 25.36 22.10 , 25.36 22.60 , 25.36 23.10 , 25.36 23.60 , 25.36 24.10 , 25.36 24.60 , 25.36
3.0 19.60 , 24.86 20.10 , 24.86 20.60 , 24.86 21.10 , 24.86 21.60 , 24.86 22.10 , 24.86 22.60 , 24.86 23.10 , 24.86 23.60 , 24.86 24.10 , 24.86 24.60 , 24.86
3.5 19.60 , 24.36 20.10 , 24.36 20.60 , 24.36 21.10 , 24.36 21.60 , 24.36 22.10 , 24.36 22.60 , 24.36 23.10 , 24.36 23.60 , 24.36 24.10 , 24.36 24.60 , 24.36
4.0 19.60 , 23.86 20.10 , 23.86 20.60 , 23.86 21.10 , 23.86 21.60 , 23.86 22.10 , 23.86 22.60 , 23.86 23.10 , 23.86 23.60 , 23.86 24.10 , 23.86 24.60 , 23.86
4.5 19.60 , 23.36 20.10 , 23.36 20.60 , 23.36 21.10 , 23.36 21.60 , 23.36 22.10 , 23.36 22.60 , 23.36 23.10 , 23.36 23.60 , 23.36 24.10 , 23.36 24.60 , 23.36
5.0 19.60 , 22.86 20.10 , 22.86 20.60 , 22.86 21.10 , 22.86 21.60 , 22.86 22.10 , 22.86 22.60 , 22.86 23.10 , 22.86 23.60 , 22.86 24.10 , 22.86 24.60 , 22.86
Note: Estimates for the identified sets for the effect $\tau_{c}$ on fiscal rules on financial deficit under Lemma 10.
Table 11: Bounds for $\tau_{uc}$ — Outcome: Deficit. $y_{min}=-762.5508$, $y_{max}=924.6157$.
$c_{2}\backslash c_{1}$ 0.0 0.5 1.0 1.5 2.0 2.5 3.0 3.5 4.0 4.5 5.0
0.0 1687.17 , 27.86 1687.17 , 27.36 1687.17 , 26.86 1687.17 , 26.36 1687.17 , 25.86 1687.17 , 25.36 1687.17 , 24.86 1687.17 , 24.36 1687.17 , 23.86 1687.17 , 23.36 1687.17 , 22.86
0.5 1687.17 , 27.86 1687.17 , 27.36 1687.17 , 26.86 1687.17 , 26.36 1687.17 , 25.86 1687.17 , 25.36 1687.17 , 24.86 1687.17 , 24.36 1687.17 , 23.86 1687.17 , 23.36 1687.17 , 22.86
1.0 1687.17 , 27.86 1687.17 , 27.36 1687.17 , 26.86 1687.17 , 26.36 1687.17 , 25.86 1687.17 , 25.36 1687.17 , 24.86 1687.17 , 24.36 1687.17 , 23.86 1687.17 , 23.36 1687.17 , 22.86
1.5 1687.17 , 27.86 1687.17 , 27.36 1687.17 , 26.86 1687.17 , 26.36 1687.17 , 25.86 1687.17 , 25.36 1687.17 , 24.86 1687.17 , 24.36 1687.17 , 23.86 1687.17 , 23.36 1687.17 , 22.86
2.0 1687.17 , 27.86 1687.17 , 27.36 1687.17 , 26.86 1687.17 , 26.36 1687.17 , 25.86 1687.17 , 25.36 1687.17 , 24.86 1687.17 , 24.36 1687.17 , 23.86 1687.17 , 23.36 1687.17 , 22.86
2.5 1687.17 , 27.86 1687.17 , 27.36 1687.17 , 26.86 1687.17 , 26.36 1687.17 , 25.86 1687.17 , 25.36 1687.17 , 24.86 1687.17 , 24.36 1687.17 , 23.86 1687.17 , 23.36 1687.17 , 22.86
3.0 1687.17 , 27.86 1687.17 , 27.36 1687.17 , 26.86 1687.17 , 26.36 1687.17 , 25.86 1687.17 , 25.36 1687.17 , 24.86 1687.17 , 24.36 1687.17 , 23.86 1687.17 , 23.36 1687.17 , 22.86
3.5 1687.17 , 27.86 1687.17 , 27.36 1687.17 , 26.86 1687.17 , 26.36 1687.17 , 25.86 1687.17 , 25.36 1687.17 , 24.86 1687.17 , 24.36 1687.17 , 23.86 1687.17 , 23.36 1687.17 , 22.86
4.0 1687.17 , 27.86 1687.17 , 27.36 1687.17 , 26.86 1687.17 , 26.36 1687.17 , 25.86 1687.17 , 25.36 1687.17 , 24.86 1687.17 , 24.36 1687.17 , 23.86 1687.17 , 23.36 1687.17 , 22.86
4.5 1687.17 , 27.86 1687.17 , 27.36 1687.17 , 26.86 1687.17 , 26.36 1687.17 , 25.86 1687.17 , 25.36 1687.17 , 24.86 1687.17 , 24.36 1687.17 , 23.86 1687.17 , 23.36 1687.17 , 22.86
5.0 1687.17 , 27.86 1687.17 , 27.36 1687.17 , 26.86 1687.17 , 26.36 1687.17 , 25.86 1687.17 , 25.36 1687.17 , 24.86 1687.17 , 24.36 1687.17 , 23.86 1687.17 , 23.36 1687.17 , 22.86
Note: Estimates for the identified sets for the effect $\tau_{uc}$ on fiscal rules on financial deficit under Lemma 10.
The results in Table 10 show that the bounds on the effects of the fiscal rules are overall uninformative, suggesting that the results are not robust to violations of the time-invariance assumptions. Tables 12 and 13, on the other hand, show that the bounds on the effects on deficit are robust both to violations on time-invariance assumptions and complementarity between treatment and confounding effects. The results in the table show that the evolution of the mean confounding effect and the mean confounded deficit at the threshold could exceed 6 euros per capita (roughly 30.000 euros given the 5.000 population rule), and the true effect of relaxing fiscal policy would still be positive. However, the setting is severely underpowered, which means that confidence intervals are uninformative regarding the true signal of the treatment effects [^3] In Tables 10 and 11, which show informative identified sets, the lower bound of the confidence interval for the lower bound is always negative..
Finally, Table 9 shows that modularity assumptions alone are not sufficient to yield informative identified sets for the treatment effects.